# Context-Parametric Inversion:Why Instruction Finetuning May Not Actually Improve Context Reliance

### v1

@@ 3.3 Key Observations
Figure 2: Accuracy on Different Knowledge Conflict Datasets We track how the model’s context-reliance (over parametric memory) evolves during instruction fine-tuning, particularly under knowledge conflicts. Counterfactual (blue) and parametric (orange) accuracy on (a) CF Quotes, (b) Biographics, and (c) World Facts versus average performance on standard benchmarks (GSM8k, MMLU, ARC, SQuAD).
Consider finetuning Llama2-7B on TULU, a general-purpose instruction tuning dataset. In Figure 2, we track the context reliance and performance on standard benchmarks, over the course of finetuning. First, observe that the average performance on standard benchmarks (GSM8k, MMLU, ARC, and SQuAD) improves with instruction finetuning as expected. Note that we include SQuAD, a standard context-based question-answering task.
On the other hand, on our question-answering datasets with counterfactual contexts, contrary to the intuition that instruction finetuning would improve dependence on user-provided context (§ 1), we observe that performance strictly decreases with instruction finetuning, after an initial expected increase. For example, on CF_World_Facts (Figure 2(c)), the context reliance initially improves from 40% to almost 90% in the initial phases of finetuning. However, it starts to decline gradually as instruction tuning progresses further. Similar observations can be made on CF_Bio dataset(Figure 2(b)). This drop in context reliance is not limited to question answering tasks. We observe a similar behavior on the CF_Quotes (Fig 2(a)), where the input instructions require an answer different from well known behaviors (Appendix A.5). On this task, the counterfactual accuracy (answering based on the input instruction) improves from 40% at zeroshot to 70%, but decreases as finetuning progresses further. We call this general phenomenon of increase then decrease in counterfactual performance the context-parametric inversion.
Context-parametric inversion appears consistently across multiple instruction tuning datasets (TULU, UltraChat, Alpaca) and model families (Llama2-7B, Pythia-6.9B, and Mistral-7B). For additional empirical results, we refer the reader to Appendix A.1. In Figure 3(b) and Appendix A.3, we also experiment with explicitly prompting the model to prioritize the context over parametric knowledge. However, the drop in context reliance persists.
@@
Our observations do not fall under the classic forgetting regime, where the performance drops monotonically on all tasks that are orthogonal (out-of-distribution) to the finetuning data. As we have shown, the performance on standard benchmarks continue to improve. Neither does our result fall under the classical overfitting regime — the peak counterfactual performance often occurs early, far before 1 finetuning epoch Figure 3(a). Further, we note that this is not simply due to memorization of related facts during instruction finetuning. In § 4.1 we show that the performance drop cannot be simply resolved by removing any overlap between facts in counterfactual datasets and the instruction finetuning datasets. In the next section, we study in detail why this phenomenon goes beyond the usual discourse on catastrophic forgetting or memorization. We perform various controlled studies to understand and isolate the cause of context-parametric inversion.
@@ 4 Why does context-parametric inversion happen?
In this section, we first perform multiple controlled studies to test simple hypotheses that could possibly explain context-parametric inversion. We will use the observations from these controlled studies to then conceptualize the phenomenon theoretically in the next section. We conduct all of our studies on the Alpaca instruction tuning dataset over Llama2-7B  unless otherwise specified.
Figure 3: (a) Counterfactual accuracy variation on CF_Quotes. Observe that the peak occurs well before the completion of even a single epoch, contrasting our observations with classical overfitting. (b) Even when explicitly prompting the model to follow the context, we observe a similar drop in context-reliance (additional results in Appendix A.3). (c) We plot the average attention score on the context for the CF_World_Facts eval set during TULU finetuning. Consistent with our theoretical analysis (§ 5), the attention on context initially rises and then falls. Note that this observation is for corroboration of our theory only and does not imply causal claims, as attention maps entangle information in deep networks.
@@ 4.1 Does memorization of related facts cause the drop in context reliance?
A straightforward explanation of the drop in context-reliance could be that models memorize additional facts seen at finetuning, and these facts bias the model’s behavior towards parametric answering for related questions. We consider filtering out any overlap of facts present in the counterfactual dataset from the finetuning data. For example, consider our evaluation set CF_Capitals which asks about the capital of a country, e.g., “What is the capital of France?” with a counterfactual context suggesting the answer as Lyon instead of Paris. Then we may filter out all examples in Alpaca that contain any country or their capital city names. This removes around $5\%$ of the training data.
Figure 4(a) compares the context-reliance of Llama2-7B finetuned on this filtered Alpaca with the standard Alpaca dataset. Interestingly, we still observe a drop in counterfactual performance after an initial increase, while specifically controlling for any train-test overlap. This highlights that context-parametric inversion is not simply because more facts gets encoded in the parametric knowledge. Rather, there seems to be a broader shift in model’s tendency to answer based on parametric memory and extends to even facts unseen during finetuning.
@@ 4.2 Lack of enough datapoints that encourage context reliance?
Another possible reason for the drop in context reliance could be that there are just not enough datapoints promoting context-reliance. Specifically, a large portion of Alpaca instruction-finetuning examples require models to assist users through pure fact recall. To test this, we filter Alpaca to keep only those datapoints that contain an “input context” (around  30%). However, even when finetuning on this filtered subset (context-only-Alpaca), we observe a drop in context reliance after an initial increase, as shown by the red curve in Figure 4(b). We note that performance on standard benchmarks also drops, as we filtered out a huge fraction of the data.
@@
Figure 4: (a) Controlling for fact overlap between train-test sets, we still observe a drop in context reliance. (b) When finetuning on context-only-Alpaca subset, a drop in context reliance is still observed. However, on a context-critical subset of Alpaca, there is no drop. (c) A drop in context reliance can also be observed when finetuning on context-based QA datasets like SQuAD.
@@ 4.3 Do all the context datapoints really need the context?
We observed above that even when finetuning only on examples with an input context, we still observe a drop in context reliance. This suggests that probably not all of these datapoints that “supposedly” have an input context really require the model to focus on context to correctly answer the question. For example, there may be datapoints where the context is aligned with the parametric knowledge that model has already seen during pretraining (Figure 1(b), non-context-critical datapoints).
To test this, we remove an additional 25% of datapoints from context-only-Alpaca that have the lowest loss, without the input context (i.e., perplexity on the target). The intuition here is that if the context aligns with the model’s parametric knowledge, the model be able to answer corresponding questions even without the context. Figure 4(b) illustrates the context reliance when fine-tuning on this filtered dataset. Notably, the context reliance does not drop in this case. Note that as expected, the performance on standard benchmarks is lower than that achieved when finetuning on full Alpaca.
The above observations indicate that the drop in context reliance during instruction finetuning is primarily driven by datapoints where the context aligns with the model’s preexisting parametric knowledge (non-context-critical datapoints). Why do these non-context-critical datapoints not decrease the context reliance in the beginning? Why does the context reliance decrease at all? In the next section, we try to answer these questions by conceptualizing this behavior theoretically in a simpler setting of one layer transformer. Later in § 6 we will use the insights from our theoretical analysis to explore mitigation strategies for this phenomenon.
@@ 5 Theoretical analysis of context-vs-parametric reliance
In the previous section (§ 4), we conducted controlled studies to isolate the cause of the drop in context reliance. We found that filtering out datapoints where the context aligns with the model’s parametric memory (§ 4.3) prevented the decrease in context reliance. Here, we try to conceptualize and understand the reason behind why non-context-critical datapoints cause a drop in context reliance.
In summary, our analysis below shows that attention to context tokens increases initially due to large gradients from data points that require context for correct predictions (context-critical points). However, as training progresses, the error on these points decreases, and gradients from the non-context-critical data points begin to dominate. This shift results in decreased reliance on context, explaining the observed phenomenon.
@@
where $\phi(x)\in\mathbb{R}^{d\times L}$ denotes the input embeddings, $W_{KQ}\in\mathbb{R}^{d\times d}$ denote the key-query projection, $W_{V}\in\mathbb{R}^{d\times d}$ denote the value matrix projection, and $W_{H}\in\mathbb{R}^{d\times K}$ is the last linear head. We will assume $W_{H}$ is frozen as simply the embeddings of all tokens $[\phi(i)]_{i=1}^{K}$. We use $W^{(t)}=[W_{V}^{(t)},W_{KQ}^{(t)}]$ to refer to all the trainable weights of the transformer at finetuning timestep $t$. We refer to instruction finetuning as Supervised Finetuning (SFT) in this section.
@@
5.1 SFT Data Composition
Our analysis hinges on the presence of two types of datapoints in the Supervised Finetuning Dataset (SFT)—(a) where context is necessary to predict the true answer given the subject and the relation (context-critical, Figure 1(b)). This can encompass both knowledge extraction based datapoints or where the context contains some instruction to for an operation to be performed on the subject. (b) where model can use either the context or its pretrained knowledge to answer. For example, the pretraining corpus $\mathcal{D}_{pre}$ may contain a set of datapoints $[s_{j},r_{j}]\in\mathcal{D}_{pre}$ $\forall$ $j\in[n_{pre}]$ that the model has already memorized (Theorem A.1,  Ghosal et al. (2024)).
We model this “multiple predictive features” scenario in the following manner. Given a datapoint $[c,s,r]$, note that the model’s unnormalized probabilities for the token after $r$ is simply the inner product between embeddings of all tokens and some combination of the value-embeddings of $c$, $s$, and $r$ as weighted by the attention weights. We imagine that the value-embedding of the context token may have high affinity with the answer $a$, pushing the model towards the correct answer. Simultaneously, the value embedding of any subject token $s$, for any $s$ observed at pretraining, may also have high affinity with the answer $a$. This allows us to categorize training points as following.
@@ 5.1 SFT Data Composition
    $\mathcal{D}_{\text{C}}$ (Context-Critical Points C): These are datapoints $([c,s,r],a)$ where the context is the only predictive feature of $a$ at SFT timestep $t=0$, in other words:
    $\left[W_{H}^{\top}W_{V}^{(0)}\phi(c)\right]_{a}>\left[W_{H}^{\top}W_{V}^{(0)}\phi(s)\right]_{a}\gg\frac{1}{\left|\mathcal{A}\right|}$ (6)
@@
    $\mathcal{D}_{\text{C+S}}$ (Non-Context-Critical Points C+S): These are datapoints $([c,s,r],a)$ where the subject-relation pair was seen during pretraining $[s,c]\in\mathcal{D}_{pre}$ and was memorized. Here, the subject is more predictive than the context of $a$ at SFT timestep $t=0$.
    $\left[W_{H}^{\top}W_{V}^{(0)}\phi(s)\right]_{a}>\left[W_{H}^{\top}W_{V}^{(0)}\phi(c)\right]_{a}\gg\frac{1}{\left|\mathcal{A}\right|}$ (7)
@@ 5.2 SFT Training Dynamic
Consider a one-layer transformer pretrained on $\mathcal{D}_{pre}$. When finetuning this transformer, with $W_{V}$ frozen, over $\mathcal{D}=$ $\mathcal{D}_{\text{C}}$$\cup$ $\mathcal{D}_{\text{C+S}}$ with ||||$\mathcal{D}_{\text{C}}$||$\geq$|| ≥ |$\mathcal{D}_{\text{C+S}}$||||, under assumptions listed in Appendix A.6, the following holds true for some learning rate $\eta^{*}$
@@
  • Second Phase At initial timestep $t=1$, the gradient of the expected loss with respect to $W_{KQ}$ observes
    $\theta_{S}^{\top}[-\nabla_{W_{KQ}}L(W^{0})]\phi(r)>0,\quad\theta_{C}^{\top}[-\nabla_{W_{KQ}}L(W^{0})]\phi(r)<0$ (10)
We defer the formal proof to Appendix A.6. Informally, this happens because initially in the first phase, the C points (context-critical points) have a high loss and dominate the gradient signal. This leads to an increase in attention weight towards the invariant context direction ($\theta_{C}$). However, as models learns to use the context, C+S points start having a comparatively larger gradient signal and push the attention back towards the invariant subject direction ($\theta_{S}$). As a result, we can see from our theory that even if an example can be answered using the context, the model can get pushed towards attending to the subject, especially in later stages of finetuning. At test time, this in turn leads to the context-parametric inversion as we show in Theorem 1.
In Figure 3(c), we infact empirically verify this, by plotting the attention score on the context, averaged over all the layers, when finetuning on the Alpaca dataset. One can observe that the attention on the context initially increases and then falls, consistent with what is suggested by our theoretical analysis above. We note here though that attention maps entangle information across the input context in deep networks. Our empirical observation here plots the attention score for the input context by approximating it to be the scores given to the corresponding positions. This is just to corroborate our theoretical insights and we do not intend to make any causal claims from this observation.
Naturally, adding pure factual recall (S points) into the training mixture exacerbates can exacerbate the shift in attention towards the subject.
@@
Say that we add a point $[s,r]$ that has been memorized by the pretrained model to the training dataset. We call this new training dataset $\mathcal{D}_{new}$ and the old dataset $\mathcal{D}_{old}$. Under assumptions listed in Appendix A.6, the gradient update with respect to $W_{KQ}$ at timestep $t=0$ observes
@@
We refer the reader to Appendix A.7 for the proof. This proposition tells us that any addition of subject points increases the attention towards the invariant subject direction $\theta_{S}$, while the attention towards the invariant context direction $\theta_{C}$ stays the same. We also saw from our theory that the value matrix places a prominent role in encoding the model’s parametric knowledge. Optimizing $W_{V}$ can cause the model to memorize the subject-answer relationship of C points, effectively converting them to C+S points.
@@
Under Assumptions in Appendix A.6, for any example $[c,s,r]\in$ $\mathcal{D}_{\text{C}}$, after the gradient step at timestep $t=0$, the value embedding of the subject token is more predictive of the label $c$.

### v3

@@ 3.3 Key Observations
Figure 2: Accuracy on Different Knowledge Conflict Datasets We track how the model’s context reliance (over parametric memory) evolves during instruction fine-tuning, particularly under knowledge conflicts. Counterfactual (blue) and parametric (orange) accuracy on (a) CF Quotes, (b) Biographics, and (c) World Facts versus average performance on standard benchmarks (GSM8k, MMLU, ARC, SQuAD).
Consider finetuning Llama2-7B on TULU, a general-purpose IFT dataset. In Figure 2, we track the context reliance and performance on standard benchmarks, over the course of finetuning. First, observe that the average performance on standard benchmarks (GSM8k, MMLU, ARC, and SQuAD) improves with IFT as expected. Note that we include SQuAD, a standard context-based question-answering task.
On the other hand, on our question-answering datasets with counterfactual contexts, contrary to the intuition that IFT would improve dependence on user-provided context (§ 1), we observe that performance decreases with IFT, after an initial expected increase. For example, on CF_World_Facts (Figure 2(c)), the context reliance initially improves from 40% to almost 90% in the initial phases of finetuning. However, it starts to decline gradually as IFT progresses further. Similar observations can be made on CF_Bio dataset (Figure 2(b)). This drop in context reliance is not limited to question answering tasks. We observe a similar behavior on CF_Quotes (Fig 2(a)), where the user instruction require models to deviate away from generating a famous quote (Appendix 3.2). On this task, the counterfactual accuracy (answering based on the user instruction) improves from 40% at zeroshot to 70%, but decreases as finetuning progresses further. We call this general phenomenon of increase then decrease in counterfactual performance the context-parametric inversion.
Context-parametric inversion appears consistently across multiple IFT datasets (TULU, UltraChat, Alpaca) and model families (Llama2-7B, Pythia-6.9B, and Mistral-7B). For additional empirical results, we refer the reader to Appendix A.1. In Appendix A.3, we also experiment with explicitly prompting the model to prioritize the context over parametric knowledge. However, the drop in context reliance persists.
@@
Our observations do not fall under the classic forgetting regime, where the performance drops monotonically on tasks that are orthogonal (out-of-distribution) to the finetuning data. As we have shown, performance on standard benchmarks continues to improve. Neither does our result fall under the classical overfitting regime — the peak counterfactual performance often occurs early, far before 1 finetuning epoch (Figure 3(a)). Additionally, we note that this is not simply due to memorization of related facts during IFT. In § 4.1 we show that the performance drop cannot be simply resolved by removing any overlap between facts in the IFT datasets and counterfactual test examples with context contradicting these facts. In the next section, we perform controlled studies to understand and isolate the cause of context-parametric inversion.
Figure 3: Not Overfitting (a) Peak performance on CF_Quotes  occurs well before the end of one epoch. (b) Attention score of LLama7B over the context for the CF_World_Facts eval set averaged over all the layers. Consistent with our theory (§ 5), the attention to context rises and falls. We do not make any causal claims from this observation about the attention dynamic in deep networks.
@@ 4 Why does context-parametric inversion happen?
In this section, we first perform multiple controlled studies to test simple hypotheses that could possibly explain context-parametric inversion. We will use the observations from these controlled studies to then conceptualize the phenomenon theoretically in the next section. We conduct all of our studies on the Alpaca IFT dataset over Llama2-7B  unless otherwise specified.
@@ 4.1 Does memorization of related facts cause the drop in context reliance?
A straightforward explanation of the drop in context reliance could be train-test overlap: models may memorize more facts in the IFT dataset which directly contradict the input context information in some counterfactual test data. This may push the model to do fact recall for these particular examples. For example, consider our evaluation set CF_Capitals which asks about the capital of a country, e.g., “What is the capital of France?” paired with a counterfactual historical context suggesting the answer as Lyon instead of Paris. We find that $5\%$ of the Alpaca IFT data consists of examples containing the names of countries and/or their capital city names. We consider filtering such examples out from the training data. Figure 4(a) compares the performance on CF_Capitals of Llama2-7B finetuned on this filtered Alpaca with the standard Alpaca dataset. Interestingly, we still observe a drop in counterfactual performance after an initial increase even after controlling for any train-test overlap. This highlights that context-parametric inversion is not simply because more facts are getting encoded in the model’s parametric knowledge during finetuning. Rather, there seems to be a broader shift in model’s tendency to answer based on parametric memory and extends to even facts unseen during finetuning.
@@ 4.2 Lack of enough datapoints that encourage context reliance?
Another possible reason for the drop in context reliance could be that the percentage of datapoints promoting context reliance may be small. Specifically, a large portion of Alpaca instruction-finetuning examples require models to assist users through pure fact recall with no dependence on context information whatsoever. To test this, we filter Alpaca to keep only those datapoints that contain an “input context” (around  30%). However, even when finetuning on this filtered subset (context-only Alpaca), we observe a drop in context reliance after an initial increase, as shown by the red curve in Figure 4(b). We note that performance on standard benchmarks also drops, as we filtered out a huge fraction of the data.
@@
Figure 4: Filtering Harmful Examples (a) Controlling for fact overlap between train-test sets, we still observe a drop in context reliance. (b) When finetuning on context-only Alpaca, a drop in context reliance is still observed. However, on a context-critical subset of Alpaca, there is no drop. (c) The drop in context reliance happens when finetuning on context-based QA datasets like SQuAD.
4.3 Context Critical vs Non-Context Critical Datapoints
Our observations from the previous section suggest that not all context-based instruction finetuning (IFT) examples effectively promote context reliance, as even when finetuning on a context-only subset of Alpaca, we observe a drop in context reliance (Figure 4(b), solid red curve). Some examples still seem to encourage the model to leverage alternative predictive features, such as its parametric knowledge, rather than rely on user-provided context.
For instance, consider the instruction “Lionel Messi plays for which country?” with the context being “Context: [overview of Messi’s career]”. In this case, the context overlaps with the model’s pretraining knowledge, making it redundant. Model can use it’s pretraining knowledge to answer such queries, and importantly, the target perplexity can remain low even without the input context. Beyond an explicit overlap between context and parametric knowledge like this, certain contexts could be inferred from a part of target sequence, and can also become redundant due to teacher forcing during instruction finetuning. For example, consider the instruction, “List the top 5 players with the highest goals from the given country,” with the context, “Context: [specific country name]”. Here the model may no longer need to focus on the context after generating the first player’s name, as the remaining answer can be inferred conditional to the previous generation. Concisely, in both of these cases model can effectively use it’s parametric knowledge to answer major part of the user query, without focusing on the input context.
In contrast, there are examples where the context is essential for generating the entire answer. Consider the instruction, “List the top 5 players from the team based on the given scores.” with the context, “Context: [Scores table]”. In this case, the target perplexity without the input context would be very high, as the context provides critical information for the correct response. Based on the above, we categorize context-based IFT examples into the following categories:
  1. (a)
    Context-Critical: The context is essential for answering the entire query and cannot be substituted with parametric knowledge or inferred from a part of the target sequence. Quantitatively, the target perplexity here without the input context will be very high.
  2. (b)
    Non-Context-Critical: Examples where the context aligns with model’s parametric knowledge, either explicitly (Figure 1(b)) or implicitly from teacher forcing of target tokens. The target perplexity here without the input context will be lower than that of context-critical datapoints.
@@ 4.4 Do all the context datapoints really need the context?
Based on the definition of context-critical and non-context-critical datapoints above, we use a simple target perplexity-based filtering to isolate a context-critical subset. We remove 25% of the datapoints from Alpaca that have the lowest target perplexity without the input context. We refer to this filtered set as “context-critical Alpaca”. Figure 4(b) (green curve) shows the context reliance when fine-tuning on this context-critical subset. Interestingly, context reliance barely drops in this case. As expected, performance on standard benchmarks is lower compared to finetuning on the full Alpaca set. We observe similar trends when fine-tuning on a context-critical subset of SQuAD, where 25% of the datapoints with the lowest target loss without context are removed. As shown in Figure 4(c) (green curve), there is no drop in context reliance when finetuning on this filtered SQuAD subset.
The above observations indicate that the drop in context reliance during IFT is primarily driven by datapoints where the context aligns with the model’s preexisting parametric knowledge (non-context-critical datapoints). Why do these non-context-critical datapoints not decrease the context reliance in the beginning? Why does the context reliance eventually decrease? In the next section, we try to answer these questions theoretically in a simpler one layer transformer setting. Later in § 6 we will use the insights from our analysis to explore mitigation strategies for this phenomenon.
@@ 5 Theoretical analysis of context-vs-parametric reliance
In the previous section (§ 4), we conducted controlled studies to isolate the cause of the drop in context reliance. We found that filtering out non-context-critical datapoints i.e. where the context is not the only predictive feature(§ 4.4) mitigated the decrease in context reliance. Here, we theoretically explain the reason behind why non-context-critical datapoints cause a drop in context reliance.
In summary, we show below that in the initial phase of finetuning, context-critical datapoints dominate the gradients, driving the model to focus on the context. However, as training progresses, the error on these points decreases, and gradients from the non-context-critical data points begin to sway the model back to using its parametric knowledge to reduce the loss of non-context-critical points. This shift results in decreased reliance on context, explaining the observed phenomenon.
@@
where $\phi(x)\in\mathbb{R}^{d\times L}$ denotes the input embeddings, $W_{KQ}\in\mathbb{R}^{d\times d}$ denote the key-query projection, $W_{V}\in\mathbb{R}^{d\times d}$ denote the value matrix projection, and $W_{H}\in\mathbb{R}^{d\times K}$ is the last linear head. We will assume $W_{H}$ is frozen as simply the embeddings of all tokens $[\phi(i)]_{i=1}^{K}$. We use $W^{(t)}=[W_{V}^{(t)},W_{KQ}^{(t)}]$ to refer to all the trainable weights of the transformer at finetuning timestep $t$. We use IFT to denote instruction finetuning in this section.
@@
5.1 IFT Data Composition
Our analysis hinges on the presence of at least two types of datapoints in the IFT dataset: (a) context-critical points, where context is the only predictive feature, given the subject and the relation (context-critical, Figure 1(b)) (b) non-context-critical points, where context is not the only predictive feature, e.g., the context overlaps with the model’s pretraining knowledge.
We assume that the pretraining corpus $\mathcal{D}_{pre}$ contains a set of datapoints $[s_{j},r_{j}]\in\mathcal{D}_{pre}$ $\forall$ $j\in[n_{pre}]$ that the model has already memorized (Theorem A.1,  Ghosal et al. (2024)). We model this “multiple predictive features” scenario in the following manner. Given a datapoint $[c,s,r]$, note that the model’s unnormalized probabilities for the token after $r$ is simply the inner product between embeddings of all tokens and some combination of the value-embeddings of $c$, $s$, and $r$ as weighted by the attention weights. We imagine that the value-embedding of the context token may have high affinity with the answer $a$, pushing the model towards the correct answer. Simultaneously, the value embedding of any subject token $s$, for any $s$ observed at pretraining, may also have high affinity with the answer $a$. This allows us to categorize training points as following.
@@ 5.1 IFT Data Composition
    $\mathcal{D}_{\text{C}}$ (Context-Critical Points C): These are datapoints $([c,s,r],a)$ where the context is the only predictive feature of $a$ at timestep $t=0$, in other words:
    $\sigma\left(W_{H}^{\top}W_{V}^{(0)}\phi(c)\right)_{a}\gg\sigma\left(W_{H}^{\top}W_{V}^{(0)}\phi(s)\right)_{a}=\frac{1}{\left|\mathcal{A}\right|}$ (6)
@@
    $\mathcal{D}_{\text{C+S}}$ (Non-Context-Critical Points C+S): These are datapoints $([c,s,r],a)$ where the subject-relation pair was seen during pretraining $[s,c]\in\mathcal{D}_{pre}$ and was memorized. Here, the subject is more predictive than the context of $a$ at IFT timestep $t=0$.
    $\sigma\left(W_{H}^{\top}W_{V}^{(0)}\phi(s)\right)_{a}>\sigma\left(W_{H}^{\top}W_{V}^{(0)}\phi(c)\right)_{a}\gg\frac{1}{\left|\mathcal{A}\right|}$ (7)
@@ 5.2 IFT Training Dynamic
Consider a one-layer transformer pretrained on $\mathcal{D}_{pre}$. When finetuning this transformer, with $W_{V}$ frozen, over $\mathcal{D}=$ $\mathcal{D}_{\text{C}}$$\cup$ $\mathcal{D}_{\text{C+S}}$ with ||||$\mathcal{D}_{\text{C}}$||$\geq$|| ≥ |$\mathcal{D}_{\text{C+S}}$||||, under assumptions listed in Appendix B.1, the following holds true for some learning rate $\eta^{*}$
@@
  • Second Phase At timestep $t=1$, the gradient of the expected loss with respect to $W_{KQ}$ observes
    $\theta_{S}^{\top}[-\nabla_{W_{KQ}}L(W^{(1)})]\phi(r)>0,\quad\theta_{C}^{\top}[-\nabla_{W_{KQ}}L(W^{(1)})]\phi(r)<0$ (10)
We defer the formal proof to Appendix B.1. Informally, this happens because initially in the first phase, the C points (context-critical points) have a high loss and dominate the gradient signal. This leads to an increase in attention weight towards the invariant context direction ($\theta_{C}$). However, as models learns to use the context, C+S points start having a comparatively larger gradient signal and push the attention back towards the invariant subject direction ($\theta_{S}$). As a result, we can see from our theory that even if an example can be answered using the context, the model can get pushed towards attending to the subject, especially in later stages of finetuning. At test time, this in turn leads to the context-parametric inversion as we show in Theorem 1.
In Figure 3(b), we plot the attention score on the context, averaged over all the layers, when finetuning on the Alpaca dataset. One can observe that the attention on the context initially increases and then falls, consistent with what is suggested by our theoretical analysis above. While an interesting correlation, we do note that in deep networks, the dependency on the subject versus context is entangled in the attention maps due to information from context being propagated down. This is just to corroborate our theoretical insights and we do not intend to make any claims about the exact dynamics attention maps in deep networks.
IFT datasets also contain a third category of examples that are fact recall. Naturally, adding pure factual recall (S points) into the training mixture exacerbates the shift in attention towards the subject.
@@
Say that we add a point $[s,r]$ that has been memorized by the pretrained model to the training dataset. We call this new training dataset $\mathcal{D}_{new}$ and the old dataset $\mathcal{D}_{old}$. Under assumptions listed in Appendix B.1, the gradient update with respect to $W_{KQ}$ at timestep $t=0$ observes
@@
We refer the reader to Appendix B.2 for the proof. This proposition tells us that any addition of subject points increases the attention towards the invariant subject direction $\theta_{S}$, while the attention towards the invariant context direction $\theta_{C}$ stays the same. Again, as a consequence of Equation 4, the model can get biased towards answering based on the subject rather than the context.
We now demonstrate that the value matrix plays a prominent role in encoding the model’s parametric knowledge. Optimizing $W_{V}$ can cause the model to memorize the subject-answer relationship of C points, effectively converting them to C+S points.
@@
Under Assumptions in Appendix B.1, for any example $[c,s,r]\in$ $\mathcal{D}_{\text{C}}$, after the gradient step at timestep $t=0$, the value embedding of the subject token is more predictive of the label $c$.